Solve 9x-3y=9, x+3y+11
Answer by tommyt3rd(5050)

9x-3y=9,
x+3y=11

adding both equations:

(9x-3y)+(x+3y)=9+11
9x +x = 20
10x=20
x=2


now using the 2nd equation

x+3y+11
2+3y=11
3y=9
y=3

solution: (2,3)



:)